1. Solutions of iron (lll) chloride, FeCl3, are used in photoengraving and to make ink. This compound can be made be the following reaction:

2F + 3Cl2 -> 2FeCl3



a. How many grams of Fe are needed to combine with 4.5 moles of Cl2?




b. If 240 g of Fe is to be used in this reaction, with adequate Cl2, how many moles of FeCl3 will be produced?


1
Expert's answer
2021-05-10T01:21:13-0400

a) V(Fe)=2/3*V(Cl2)=4.5mol*2/3=3mol

M(Fe)=56g/mol

m(Fe)=3mol*56g/mol=168g


b) V(Fe)=240g/56g/mol=4.29mol

V(Fe)=V(FeCl3)=4.29mol